Journalist Report – March 03rd

Crew 206 – 03/03/2019

Sol 7

Author: Benjamin Auzou, Journalist

"Martian Sol-Off"

For our first day-off of the mission, we played cards, begin to complete the Mars puzzle of the MDRS. But above all we spend a lot of time in the kitchen: Gaspard and Aurélien prepared pancakes for breakfast, Jérémy and Cerise baked French baguettes, and Aurélien and I cooked a gratin dauphinois (potatoes cooked in cream and cheese baked in the oven). All that was bettered by the basil leafs brought by Norbert from the Green-Hab.

Outside the Hab, the weather changed all along the day. After a sunny morning and a rainy noon, the afternoon was shared between grey clouds and hail. Even if it wasn’t Mars-like, the sound of rain and hail on the Habs roof was relaxing. In the same time our furious Monopoly Deal games were the great activity of the afternoon, even if Gaspard’s luck with cards drawing killed the competition.

This day-off is for us the occasion to present you the results of our water monitoring. We are counting the water we use in different categories: drinking, cooking, dishes, hygiene, flushes and Green-Hab. Thanks to that we managed to reduce our consumption and realized a precise monitoring. We measured in average 88% of our water consumption and classified that in the categories for each sol spend in the MDRS.

Today we recharged our batteries for the week to come that will be busy between the first tests our own simulation spacesuit and the follow-up of the human factors experiments.

More than one third of our mission is done, this first week was balanced between the first marks in the station, the first scientific experiments but also the first difficulties. As in all space missions, we are facing difficulties and we are led to adapt ourselves and react with what we have here in the station. More than ever, we are feeling that we make science evolve in the right way to Mars.

Operations Report – March 03rd

Crew 206 Operations Report 03 – MAR – 2019

SOL: 7

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): 15 hours

Solar— SOC % – Turned off Gen : 100%, Turned on Gen : 80%

Diesel Reading – 75%

Propane Reading – 81%

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – 450 gallons

Water differential (static tank) – 10 gallons

Gallons transferred: Not Applicable

Water in GreenHab – 254 gallons

Static to Loft Pump used – Yes

Water Meter: 0141642,1 gallons

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not used

Hours: 53,0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not used

Hours: 77,3 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 78,4 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Nothing to report.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Not used.

Staff Crew Car used and why, where? Off Campus

General notes and comments: Nothing to report.

Summary of internet: Still issues today. We’re beginning to investigate local devices, I’ll keep working on it with BDM CapCom.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Some rain and hail again today, we prevented major leaks from doing water damage.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report.

GreenHab Report – March 03rd

Crew 206 Green Hab Report – 03-MAR-2019

GreenHab Officer: Norbert Pouzin

Environmental control: Heating and Cooling w/ ambient air (3 hrs)

Shade Cloth (40%)

Average temperatures:

Low: 19.0°C

High: 28.8°C

Hours of supplemental light: 0 Hours

Daily water usage for crops: 3.2 gallons

Water in Blue Tank: 254.2 gallons

Crops watered at 0900 hours (1.05 gallons) and 1800 hours (2.15 gallons):


1) 09.00 AM 24.9° C / 29 %

2) 12.30 AM 26.4° C / 26 %

3) 16:00 PM 22.3° C / 33 %

4) 18:00 PM 23.7° C / 36 %

Changes to crops: None.

Narrative: Today was a rest day, so we didn’t work on the greenhab. We only harvested some herbs for our delicious meals.

Harvest: 8 g of basil and 2 g of parsley.

Support/supplies needed: None.

Sol Summary – March 03rd

Crew 206 March 3rd 2019

Sol 7

Summary Title: Monopoly afternoon

Author: Jeremy Auclair

Mission Status: All nominal

Sol Activity Summary: This morning there was no wake up alarms, but we all got up before 09:00 am. We ate pancakes and started looking for some fun games to play. The best one we found was Monopoly Deal, it’s simple, fast and fun, and we loved it. While some of us started working on the Mars puzzle in the morning, the rest of us prepared some bread and a "gratin dauphinois" (potatoes cooked in cream and cheese, baked in the oven). We ate really well, great for crew morale! The bread also looks great, we kept it for dinner. We played Monopoly a good part of the afternoon, Cerise made some good progress on the puzzle, and after that we read a bit and worked on our reports. To sum up it was a calm day which we really enjoyed!

Look ahead plan: Tomorrow we hope we will be able to go on EVA to finally test the prototype space suit. I also hope that I will make some progress on TELEOP.

Anomalies in work: None

Weather: Sunny in the morning, cloudy and rainy in the afternoon.

Crew physical status: Fine

EVA: None

Reports to be filed: GreenHab, Journalist, Operation

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site